A grandparent is eight times the age of a child. One year from now the grandparent will be seven times the age the child will be then. Find the present age of each
(I posted another problem that I put under the wrong word problem topic. I am sorry about that)

Answer by ikleyn(52803)

.
G     = 8*C        (1)    ("A grandparent is eight times the age of a child.")

G + 1 = 7*(C+1)    (2)    ("One year from now the grandparent will be seven times the age the child will be then.")



From (1), substitute the expression of G into equation (2).  You will get


8C + 1 = 7*(C+1),


8C + 1 = 7C + 7  ====>  8C - 7C = 7 - 1  ====>  C = 6.


Answer.  The child is 6 years old.  The grandfather/grandmother is 6*8 = 48 years old.

Solved.